If y varies directly as x, and y = 7 when x =5, find y when x =10
y= blank?

Answer by Alan3354(69443)

If y varies directly as x, and y = 7 when x =5, find y when x =10
---------------
y = k*x
7 = k*5
k = 7/5 = 1.4
------
@ x = 10:
y = 1.4*10
y = 14
====================
Or using proportions:
y = k*x
y/x = k (a constant)
7/5 = k = y/10
7/5 = y/10
5y = 70
y = 14